OpenStudy (anonymous):

A box has a base with an area of 55 cm2 and a height of h. Which equation can be used to find V, the volume of the box? V = 55 × 55 × h V = 55/h V = 55 + h V = 55h

OpenStudy (anonymous):

There are 4 sides of the box so the total surface area of the box is x^2+4xh and that is ... Solve that equation for h: x^2+4xh = 1000 h = (1000-x^2)/4x so the Volume ... the volume function to find the maximum volume: V=250(18.26)-(18.26)^3/4 V= ... in a right circular cone of height 3.5 and radius (at the base) equal to 5.5